OpenStudy (anonymous):

Write the equation of a line that is parallel to the line 2x+y=8 and passes through the point (-2,1)

OpenStudy (compassionate):

Parallel lines have the same slope. So first, find out the slope of 2x + y = 8, by putting it in y = mx + b

OpenStudy (anonymous):

y=2x+8

OpenStudy (compassionate):

So the slope is 2 Now we're given the points (-2, 1) So slope = 2 Points = (-2, 1) Plug those into: y - y1 = m(x - x1) Where y1 and x1 are points and m is the slope

OpenStudy (anonymous):

So the equation is y-1=2x+4

OpenStudy (compassionate):

No put that in y = mx + b

OpenStudy (anonymous):

y=2x+4

OpenStudy (compassionate):

There's your equation.

OpenStudy (anonymous):

Thanks

OpenStudy (compassionate):

---------------> Best responses are appreciated.